(a) You and your new alien friend, Howard need to escape area 51 fast! Luckily, Howard knows of a high-end sports car you can use in your escape. Assuming you don't get shot in the process, what is the minimum acceleration necessary to reach a linear displacement of 5.00 km in under 2.00 minutes? Assume you start from rest and maintain a constant acceleration. (b) Howard's Bugatti has a maximum acceleration around 11.0 m/s2 and a top speed of about 260.0 mph. What is the least amount of time it will take you to reach a linear displacement of 5.00 km if you accelerate at a constant rate of 11.0 m/s until reaching top speed and then cruise at the top speed for the remainder of the trip into Mexico? (
1
Expert's answer
2019-09-23T09:34:01-0400

a)Using the law for uniformly accelerated motion we can express a travelled distance in time "t" (and we use the condition that we start from the rest "{v_0} = 0")


"l = {v_0}t + \frac{{a{t^2}}}{2} = \frac{{a{t^2}}}{2}"


Expressing acceleration


"a = 2\frac{l}{{{t^2}}}"


We get minimal value of acceleration needed to reach the given displacement if we give a travelled distance is equal to the given displacement. Now do the calculations


"{a_{\min }} = 2\frac{l}{{{t^2}}} = 2\frac{{5[{\text{km}}]}}{{{2^2}[{{\min }^2}]}} = 2\frac{{5 \cdot {{10}^3}[{\text{m}}]}}{{{2^2} \cdot {{60}^2}[{{\text{s}}^2}]}} = \frac{{25}}{{36}}[\frac{{\text{m}}}{{{{\text{s}}^2}}}] \approx 0.7[\frac{{\text{m}}}{{{{\text{s}}^2}}}]"

b)At the first let's evaluate the time needed to reach the top speed from the rest (denote it by "\tau") . Use the law for uniformly accelerated motion


"v = {v_0} + a\tau = a\tau""\tau= \frac{v}{a}"

We will need to convert mph in m/s, it's can be done by multiplying by "{0.447}". Thus


"\tau = \frac{{{v_{\max }}}}{a} = \frac{{260[{\text{mph]}}}}{{11[\frac{{\text{m}}}{{{{\text{s}}^{\text{2}}}}}]}} \approx \frac{{0.447 \cdot 260[\frac{{\text{m}}}{{\text{s}}}{\text{]}}}}{{11[\frac{{\text{m}}}{{{{\text{s}}^{\text{2}}}}}]}} \approx 10.566[{\text{s}}]"

In this time the car travelled


"d = \frac{{a{t^2}}}{2} = \frac{1}{2} \cdot 11[\frac{{\text{m}}}{{{{\text{s}}^{\text{2}}}}}] \cdot {10.566^2}[{{\text{s}}^2}] = 614.022[{\text{m}}]"

The rest part of the full distance "x = l - d" the car will move with constant (top) speed. It will take the time


"t = \frac{{l - d}}{{{v_{\max }}}} = \frac{{5[{\text{km}}] - 614.022[{\text{m}}]}}{{260[{\text{mph]}}}} \approx \frac{{5 \cdot {{10}^3}[{\text{m}}] - 614.022[{\text{m}}]}}{{0.447 \cdot 260[\frac{{\text{m}}}{{\text{s}}}{\text{]}}}} \approx 37.739[{\text{s}}]"

Thus the full time to reach the given displacement is


"{t_{{\text{full}}}} = t + \tau = 10.566[{\text{s}}] + 37.739[{\text{s}}] = 48.305[{\text{s}}] \approx 48.3[{\text{s}}]"

This is the answer to part b)